Bill Zimmerman is evaluating two new business opportunities. Each of the opportunities shown below has a ten-year life. Bill uses a 11% discount rate.

Option 1 Option 2
Equipment purchase and installation $72,000 $83,270
Annual cash flow $28,800 $30,860
Equipment overhaul in year 3 $4,620 -
Equipment overhaul in year 5 - $5,990


Calculate the net present value of the two opportunities. (Round present value factor calculations to 4 decimal places, e.g. 1.2514 and the final answers to 0 decimal places, e.g. 59,991.)

Option 1 Option 2
Net present value


Calculate the profitability index of the two opportunities. (Round answers to 2 decimal places, e.g. 15.25.)

Option 1 Option 2
Profitability Index

Which option should Bill choose?
Answer #1

Answer- 1)- Net Present Value- Option 1 =$94231

Option 2 =$94916

Explanation-

Bill Zimmerman
Net Present Value (Option 1)
Particulars Cash Flows Present Value Factor @11% Present value
(a) (b) (c=a*b)
$ $ $
Annual Cash flow (For 10 years) 28800 5.8892 169609
Equipment purchase & installation -72000 1 -72000
Equipment overhaul in year 3 -4620 0.7312 -3378
Net Present Value 94231
Bill Zimmerman
Net Present Value (Option 2)
Particulars Cash Flows Present Value Factor @11% Present value
(a) (b) (c=a*b)
$ $ $
Annual Cash flow (For 10 years) 30860 5.8892 181741
Equipment purchase & installation -83270 1 -83270
Equipment overhaul in year 5 -5990 0.5935 -3555
Net Present Value 94916

2)- Profitablity Index- Option 1 = 1.31

Option 2= 1.14

Explanation-

Calulation of Profitablity Index
Particulars Option
1 2
Net Present Value $ (A) 94231 94916
Investment required $ (B) 72000 83270
Profitablity Index C=A/B 1.31 1.14

Bill should choose option 1.